# From Human Hands to Robot Arms: Manipulation Skills Transfer via Trajectory Alignment

### v1

@@ Abstract
Learning diverse manipulation skills for real-world robots is severely bottlenecked by the reliance on costly and hard-to-scale teleoperated demonstrations. While human videos offer a scalable alternative, effectively transferring manipulation knowledge is fundamentally hindered by the significant morphological gap between human and robotic embodiments. To address this challenge and facilitate skill transfer from human to robot, we introduce Traj2Action,a novel framework that bridges this embodiment gap by using the 3D trajectory of the operational endpoint as a unified intermediate representation, and then transfers the manipulation knowledge embedded in this trajectory to the robot’s actions. Our policy first learns to generate a coarse trajectory, which forms an high-level motion plan by leveraging both human and robot data. This plan then conditions the synthesis of precise, robot-specific actions (e.g., orientation and gripper state) within a co-denoising framework. Extensive real-world experiments on a Franka robot demonstrate that Traj2Action boosts the performance by up to 27% and 22.25% over $\pi_{0}$ baseline on short- and long-horizon real-world tasks, and achieves significant gains as human data scales in robot policy learning. Our project website, featuring code and video demonstrations, is available at https://anonymous.4open.science/w/Traj2Action-4A45/.
1 Introduction
@@
Enabling robots to master a diverse array of manipulation skills in the real world presents a formidable challenge, primarily due to the data-hungry nature of modern policy learning (Kim et al., 2024; Black et al., ; Team et al., 2024). The prevailing paradigm of imitation learning relies on large-scale datasets of expert demonstrations (O’Neill et al., 2024; Khazatsky et al., 2024), which are typically gathered through time-consuming teleoperation (Wu et al., 2023; Fu et al., 2024; Zhao et al., 2023; Orbik, 2021). This process not only requires significant investment in specialized hardware but also demands extensive operator training, rendering it a critical bottleneck for scaling up robotic capabilities.
@@
While human videos offer a cost-effective and abundant alternative data source, their direct application is fundamentally hindered by the significant morphological gap between embodiments, as they lack the robot-specific action labels required for direct mimicry. Prior works have attempted to learn a shared visual representation across both human and robot videos (Wang et al., 2023), formulating reward functions (Ma et al., 2022; Shao et al., 2021; Ma et al., 2023), or using human demonstrations as high-level context for robot action prediction (Zhu et al., 2025; Shah et al., 2025). However, these indirect approaches often fail to leverage the rich, explicit motion signals embedded within human actions. More direct lines of research attempt to define a unified action space or to transfer motion skills in a way that can be effectively utilized by robots. Yet, these strategies often introduce their own limitations. Some methods adopt a two-stage pipeline Bharadhwaj et al. (2024); Bi et al. (2025); Luo et al. (2025), where models are first trained on human data and then adapted with robot data, but this sequential procedure can constrain the final quality of robot policy. Other joint-training approaches (Yang et al., 2025; Qiu et al., 2025) are typically confined to kinematically similar embodiments like humanoid robots, while methods (Ren et al., 2025; Kareer et al., 2025) that directly align mismatched poses (e.g., human hand to parallel gripper) struggle to establish a physically meaningful correspondence. Even though some works (Bharadhwaj et al., 2024) adopt simplified representations like rigid transformations to mitigate the embodiment gap, fundamental structural differences mean that the physical interpretation of rotations remains misaligned. Furthermore, other methods (Park et al., 2025; Liu et al., 2025) impose heavy constraints, either requiring strictly paired human-robot demonstrations or relying on complex motion retargeting pipelines, which restricts their applicability and negates the cost-benefit of using human data.
@@
To address these limitations, we propose Traj2Action, a framework that transfers motion knowledge from human to robot learning process by leveraging 3D trajectories as a robust intermediate representation. Our method follows a coarse-to-fine action prediction paradigm for robot, where coarse trajectory planning derived from both human and robot demonstrations guides fine-grained robot action learning. Specifically, we introduce a unified trajectory representation of human hand and robot end-effector 3D positions, which reduces embodiment discrepancies and allows abundant, cost-effective human data to improve coarse trajectory planning. We then utilize a co-denoising training scheme over trajectories and actions, which enables the policy to focus on generating robot actions with precise orientation and gripper states under the guidance of the coarse trajectory plan. To further bridge the observational gap, we designed a hand wrist-mounted camera to supply ego-view for human data collection, which mimics the robot’s ego-centric view observation and thus enhances cross-embodiment training effectiveness and consistency.
@@
By leveraging human demonstrations, our method surpass a traditional VLA baseline on short- and long-horizon real-world tasks by a large margin. Besides, as human data scales, our method shows a significant performance gain in robot manipulation tasks. Furthermore, we show that our method allows for replacing a significant amount of expensive robot data with low-cost human demonstrations while achieving comparable final policy performance, validating the effectiveness of using low-cost human demonstration data for robot learning.
2 Related Works
@@
Robotic Imitation Learning. Imitation learning (IL) (Schaal, 1996; Pomerleau, 1988; Atkeson & Schaal, 1997) has emerged as a dominant paradigm for tackling complex robotic manipulation tasks, enabling agents to learn expert behaviors from visual observations and language instructions. Recent years have witnessed remarkable progress in this domain, largely propelled by sophisticated policy architectures and novel training schemes like Action Chunking Transformer (ACT) (Zhao et al., 2023), Diffusion Policy series (Chi et al., 2023; Reuss et al., 2024). These policies excel at mapping high-dimensional sensory inputs and proprioceptive states to robot actions.
@@
Despite their success, the performance of these state-of-the-art methods is fundamentally contingent upon access to large-scale, expert-level robot demonstration datasets (Khazatsky et al., 2024; O’Neill et al., 2024). This dependency has co-evolved with the rise of generalist Vision-Language-Action Models (Black et al., ; Kim et al., 2024), which are trained on these massive datasets with the goal of creating a single, generalizable policy that spans diverse tasks and environments. While significant efforts have been made to streamline the burdensome process of teleoperated data collection—through innovations such as kinesthetic teaching (Wu et al., 2023) and VR-based mirroring (Ding et al., 2024; Shaw et al., 2024; Park & Agrawal, 2024)—a critical bottleneck persists. Expanding these datasets to cover a long tail of diverse tasks and environments remains prohibitively challenging and expensive. This fundamental limitation motivates the search for alternative paradigms capable of learning effective policies from more abundant, scalable, and lower-cost data sources.
@@
Learning from Human Demonstration. Given the challenges associated with collecting teleoperated robot data, videos of human activity have become a popular alternative data source due to their vast availability and ease of collection. One line of work utilizes large-scale human video datasets for pre-training powerful visual representations (Nair et al., 2022; Karamcheti et al., 2023; Baker et al., 2022). However, they lack the direct training of action labels from human videos and hinders transferability to robot learning.
@@
Another line of research (Bi et al., 2025; Luo et al., 2025) attempt to directly pre-train a robot policy on human videos by using hand poses as action labels, followed by lightweight fine-tuning on the target robot. However, this strategy’s fail to address the significant morphological gap between a human hand and a robotic end-effector, which hinders effective skill transfer. Several approaches (Wang et al., 2023; Zakka et al., 2022; Xu et al., 2023) focus on task-specific adaptation, where human videos are used as a reference to enable few-shot learning on new tasks with a small number of robot demonstrations. However, these approaches requires strictly paired human-robot datasets, a constraint that significantly increases data collection difficulty and cost.
@@
To specifically address the morphological gap, researchers have proposed several strategies. For kinematically similar embodiments, such as dexterous hands, a common approach is to model both the human and robot hand with a unified 3D mesh and use motion retargeting to translate human keypoints into robot joint positions (Yang et al., 2025; Park et al., 2025; Qiu et al., 2025; Luo et al., 2025). For more heterogeneous pairings, like a human hand and a simple parallel gripper, methods often build a unified representation in an intermediate space. This includes predicting the future pixel locations of 2D keypoint tracks in the image space (Ren et al., 2025) or jointly regressing disparate human and robot end-effector poses (Qiu et al., 2025).
@@
3 Method
In this section, we present Traj2Action for transfering manipulation skills from human demonstrations to robot. Our methodology is structured to answer two key questions: (1) How can we unify human and robot demonstrations despite their significant embodiment differences? and (2) How can this unified knowledge be effectively integrated into the robot’s learning process? To this end, we introduce unification strategy for human and robot, which involves representing motions in a unified trajectory space (Section 3.1). We then detail the Traj2Action, which is designed to translate this unified trajectory representation into fine-grained robot actions (Section 3.2). Data collection pipeline for acquiring the necessary cross-embodiment demonstrations is introduced finally (Section 3.3).
3.1 Unified Trajectory Space
@@ 3.2 Traj2Action
$\mathcal{L}_{\text{traj}}(\mathcal{T})=\mathbb{E}_{\tau,\mathbf{z},\mathbf{T}^{*}}\left[\|g_{\mathcal{T}}(\mathbf{T}_{\tau},\tau,\mathbf{I}_{t},\mathcal{P},\mathbf{q}^{\prime}_{t})-(\mathbf{z}-\mathbf{T}^{*}_{t+1:t+H})\|^{2}\right],$ (1)
@@
The Action Expert is trained to denoise a noisy action $\mathbf{a}_{\tau}=\tau\cdot\mathbf{a}^{*}_{t+1:t+H}+(1-\tau)\mathbf{z}$, conditioned on the noisy action $\mathbf{a}_{\tau}$ and noisy trajectory $\mathbf{T}_{\tau}$ with a causal attention pattern as shown in the right of Figure. 1. This conditioning is the key mechanism for integrating the manipulation prior from the corresponding denoising process of unified trajectory. The action denoising loss is calculated by:
$\mathcal{L}_{\text{action}}(\theta)=\mathbb{E}_{\tau,\mathbf{z},\mathbf{a}^{*}}\left[\|\pi_{\theta}(\mathbf{a}_{\tau},\mathbf{T}_{\tau},\tau,\mathbf{I}_{t},\mathcal{P},\mathbf{q}_{t})-(\mathbf{z}-\mathbf{a}^{*}_{t+1:t+H})\|^{2}\right],$ (2)
@@
3.3 Data Collection for Human and Robot
@@
Human Hand Motion Capture System. We capture high-fidelity 3D human hand poses using a calibrated multi-camera system (as shown in Figure 2). Our pipeline uses Google MediaPipe (Zhang et al., 2020) to detect 2D hand keypoints from three views, then fits the parametric MANO (Romero et al., 2017) 3D hand model to these keypoints by minimizing reprojection error. This yields an accurate 3D hand motion, from which we extract the operational endpoint’s trajectory.
@@
Robot Data Collection. Our platform features a Franka Research 3 arm with a UMI Gripper (Chi et al., 2024). An expert teleoperates the robot via a SpaceMouse. We record proprioceptive data and images from a static third-person camera and a wrist-mounted camera. The robot’s end-effector 3D position is directly recorded as its trajectory. Simultaneously, the end-effector’s 3D orientation and the gripper state are also recorded. Combined with the 3D position, these constitute the full action label for policy training. To enhance policy robustness, we randomize the robot’s starting pose for each demonstration, which improves the policy’s ability to recover from errors. For more details, please refer to Appendix A.2.
4 Experiments
Real-World Tasks. As shown in Figure 3, four distinct real-world tasks are designed to evaluate the policy’s capabilities across basic manipulation, language grounding, long-horizon planning, and precise control. The tasks are detailed as follows:
Task 1: pick up the water bottle (PWB): The robot is tasked with locating a water bottle placed on a tabletop, moving towards it, and grasping it successfully. This task evaluates the model’s fundamental pick-and-place capabilities. Task 2: pick up the tomato and put it in the yellow/blue tray (PTT): The workspace contains a tomato and two trays, one yellow and one blue. The robot must pick up the tomato and place it into the tray specified by a language command (e.g., “the yellow tray”). This task tests the policy’s ability to task instructions to specific objects and goals. Task 3: stack the rings on the pillar (SRP): The scene includes a pillar (composed of a yellow column and a blue base), a yellow ring, and a red ring. The robot needs to pick up both rings, one by one, and place them onto the pillar. This task assesses multi-step object manipulation and precision location, which demands that the policy has precise action control. Task 4: stack the paper cups (SPC): Three paper cups are placed on the table. The robot is required to stack them sequentially to form a single tower. This task evaluates the policy’s ability to handle deformable objects and perform iterative, precise placement. We categorize these tasks as either Short-Horizon or Long-Horizon. Tasks 1 and 2 are Short-Horizon, evaluating fundamental skills like object grasping and language grounding. In contrast, Tasks 3 and 4 are significantly more demanding Long-Horizon tasks that test the policy’s planning capability. They not only require the aforementioned skills but also critically assess the policy’s ability to perform high-level planning for complex, sequential operations.

### v2

# Traj2Action: A Co-Denoising Framework for Trajectory-Guided Human-to-Robot Skill Transfer

@@ Abstract
Learning diverse manipulation skills for real-world robots is severely bottlenecked by the reliance on costly and hard-to-scale teleoperated demonstrations. While human videos offer a scalable alternative, effectively transferring manipulation knowledge is fundamentally hindered by the significant morphological gap between human and robotic embodiments. To address this challenge and facilitate skill transfer from human to robot, we introduce Traj2Action, a novel framework that bridges this embodiment gap by using the 3D trajectory of the operational endpoint as a unified intermediate representation, and then transfers the manipulation knowledge embedded in this trajectory to the robot’s actions. Our policy first learns to generate a coarse trajectory, which forms a high-level motion plan by leveraging both human and robot data. This plan then conditions the synthesis of precise, robot-specific actions (e.g., orientation and gripper state) within a co-denoising framework. Our work centers on two core objectives: first, the systematic verification of the Traj2Action framework’s effectiveness—spanning architectural design, cross-task generalization, and data efficiency and second, the revelation of key laws that govern robot policy learning during the integration of human hand demonstration data. This research focus enables us to provide a scalable paradigm tailored to address human-to-robot skill transfer across morphological gaps. Extensive real-world experiments on a Franka robot demonstrate that Traj2Action boosts the performance by up to 27% and 22.25% over $\pi_{0}$ baseline on short- and long-horizon real-world tasks, and achieves significant gains as human data scales in robot policy learning.
I Introduction
@@
Enabling robots to master a diverse array of manipulation skills in the real world presents a formidable challenge, primarily due to the data-hungry nature of modern policy learning [16, 6, 37]. The prevailing paradigm of imitation learning relies on large-scale datasets of expert demonstrations [23, 15], which are typically gathered through time-consuming teleoperation [39, 12, 44, 24]. This process not only requires significant investment in specialized hardware but also demands extensive operator training, rendering it a critical bottleneck for scaling up robotic capabilities.
@@
While human videos offer a cost-effective and abundant alternative data source, their direct application is fundamentally hindered by the significant morphological gap between embodiments, as they lack the robot-specific action labels required for direct mimicry. Prior works have attempted to learn a shared visual representation across both human and robot videos [38], formulating reward functions [20, 35, 19], or using human demonstrations as high-level context for robot action prediction [45, 34]. However, these indirect approaches often fail to leverage the rich, explicit motion signals embedded within human actions. More direct lines of research attempt to define a unified action space or to transfer motion skills in a way that can be effectively utilized by robots. Yet, these strategies often introduce their own limitations. Some methods adopt a two-stage pipeline [4, 5, 18], where models are first trained on human data and then adapted with robot data, but this sequential procedure can constrain the final quality of robot policy. Other joint-training approaches [41, 29] are typically confined to kinematically similar embodiments like humanoid robots, while methods [30, 14] that directly align mismatched poses (e.g., human hand to parallel gripper) struggle to establish a physically meaningful correspondence. Even though some works [4] adopt simplified representations like rigid transformations to mitigate the embodiment gap, fundamental structural differences mean that the physical interpretation of rotations remains misaligned. Furthermore, other methods [25, 17] impose heavy constraints, either requiring strictly paired human-robot demonstrations or relying on complex motion retargeting pipelines, which restricts their applicability and negates the cost-benefit of using human data.
@@
To address these limitations, we propose Traj2Action, a framework that transfers motion knowledge from human to robot learning process by leveraging 3D trajectories as a robust intermediate representation. Our method follows a coarse-to-fine action prediction paradigm for robot, where coarse trajectory planning derived from both human and robot demonstrations guides fine-grained robot action learning. Specifically, we introduce a unified trajectory representation of human hand and robot end-effector 3D positions, which reduces embodiment discrepancies and allows abundant, cost-effective human data to improve coarse trajectory planning. Importantly, we do not claim that trajectory alone suffices to solve robotic manipulation; rather, it acts as a principled abstract mediator bridging high-level planning and low-level execution across distinct embodiments. We then utilize a co-denoising training scheme over trajectories and actions, which enables the policy to focus on generating robot actions with precise orientation and gripper states under the guidance of the coarse trajectory plan. To further bridge the observational gap, we designed a hand wrist-mounted camera to supply ego-view for human data collection, which mimics the robot’s ego-centric view observation and thus enhances cross-embodiment training effectiveness and consistency.
@@
By leveraging human demonstrations, our method surpass a traditional VLA baseline on short- and long-horizon real-world tasks by a large margin. Besides, as human data scales, our method shows a significant performance gain in robot manipulation tasks. Furthermore, we show that our method allows for replacing a significant amount of expensive robot data with low-cost human demonstrations while achieving comparable final policy performance, validating the effectiveness of using low-cost human demonstration data for robot learning. Our core goal is to first validate the effectiveness of trajectory-guided human-to-robot skill transfer with high-fidelity human demonstration data, then extend this paradigm to large-scale low-cost internet video datasets in future work.
II Related Works
@@
Robotic Imitation Learning. Imitation learning (IL) [33, 27, 1] has emerged as a dominant paradigm for tackling complex robotic manipulation tasks, enabling agents to learn expert behaviors from visual observations and language instructions. Recent years have witnessed remarkable progress in this domain, largely propelled by sophisticated policy architectures and novel training schemes like Action Chunking Transformer (ACT) [44], Diffusion Policy series [9, 31]. These policies excel at mapping high-dimensional sensory inputs and proprioceptive states to robot actions.
@@
Despite their success, the performance of these state-of-the-art methods is fundamentally contingent upon access to large-scale, expert-level robot demonstration datasets [15, 23]. This dependency has co-evolved with the rise of generalist Vision-Language-Action Models [6, 16], which are trained on these massive datasets with the goal of creating a single, generalizable policy that spans diverse tasks and environments. To address this data dependency, some works [21] have explored leveraging simulated data to compensate for the scarcity of real-world data. While significant efforts have been made to streamline the burdensome process of teleoperated data collection through innovations such as kinesthetic teaching [39] and VR-based mirroring [11, 36, 26], a critical bottleneck persists. Expanding these datasets to cover a long tail of diverse tasks and environments remains prohibitively challenging and expensive. This fundamental limitation motivates the search for alternative paradigms capable of learning effective policies from more abundant, scalable, and lower-cost data sources.
@@
Learning from Human Demonstration. Given the challenges associated with collecting teleoperated robot data, videos of human activity have become a popular alternative data source due to their vast availability and ease of collection. One line of work utilizes large-scale human video datasets for pre-training powerful visual representations [22, 13, 3]. However, they lack the direct training of action labels from human videos and hinders transferability to robot learning.
@@
Another line of research [5, 18] attempt to directly pre-train a robot policy on human videos by using hand poses as action labels, followed by lightweight fine-tuning on the target robot. However, this strategy fail to address the significant morphological gap between a human hand and a robotic end-effector, which hinders effective skill transfer. Several approaches [38, 42, 40] focus on task-specific adaptation, where human videos are used as a reference to enable few-shot learning on new tasks with a small number of robot demonstrations, or map human manipulation actions to robot actions to provide a prior [2] or reference [8]. However, these approaches requires strictly paired human-robot datasets, a constraint that significantly increases data collection difficulty and cost.
@@
To specifically address the morphological gap, researchers have proposed several strategies. For kinematically similar embodiments, such as dexterous hands, a common approach is to model both the human and robot hand with a unified 3D mesh and use motion retargeting to translate human keypoints into robot joint positions [41, 25, 29, 18]. For more heterogeneous pairings, like a human hand and a simple parallel gripper, methods often build a unified representation in an intermediate space. This includes predicting the future pixel locations of 2D keypoint tracks in the image space [30] or jointly regressing disparate human and robot end-effector poses [29]. Methods such as AnyTeleop [28] rely on optimization-based hand pose retargeting to align human hand poses with dexterous robot hand poses. This optimization-based retargeting logic dependent on manual configuration is unfavorable for the expansion of large-scale datasets in vision-based robot teleoperation; thus, finding a simple and unified representation space is crucial for large-scale data collection in the field of robot training.
@@
III Method
In this section, we present Traj2Action for transfering manipulation skills from human demonstrations to robot. Our methodology is structured to answer two key questions: (1) How can we unify human and robot demonstrations despite their significant embodiment differences? and (2) How can this unified knowledge be effectively integrated into the robot’s learning process? To this end, we introduce unification strategy for human and robot, which involves representing motions in a unified trajectory space (Section III-A). We then detail the Traj2Action, which is designed to translate this unified trajectory representation into fine-grained robot actions (Section III-B). Data collection pipeline for acquiring the necessary cross-embodiment demonstrations is introduced finally (Section III-C).
III-A Unified Trajectory Space
@@ III-B Traj2Action
$\begin{split}\mathcal{L}_{\text{traj}}(\mathcal{T})=\mathbb{E}_{\tau,\mathbf{z},\mathbf{T}^{*}}\bigg[&\|g_{\mathcal{T}}(\mathbf{T}_{\tau},\tau,\mathbf{I}_{t},\mathcal{P},\mathbf{q}^{\prime}_{t})\\ &-(\mathbf{z}-\mathbf{T}^{*}_{t+1:t+H})\|^{2}\bigg],\end{split}$ (1)
@@
The Action Expert is trained to denoise a noisy action $\mathbf{a}_{\tau}=\tau\cdot\mathbf{a}^{*}_{t+1:t+H}+(1-\tau)\mathbf{z}$, conditioned on the noisy action $\mathbf{a}_{\tau}$ and noisy trajectory $\mathbf{T}_{\tau}$ with a causal attention pattern as shown in the right of Figure 1. This conditioning is the key mechanism for integrating the manipulation prior from the corresponding denoising process of unified trajectory. The action denoising loss is calculated by:
$\begin{split}\mathcal{L}_{\text{action}}(\theta)=\mathbb{E}_{\tau,\mathbf{z},\mathbf{a}^{*}}\bigg[&\|\pi_{\theta}(\mathbf{a}_{\tau},\mathbf{T}_{\tau},\tau,\mathbf{I}_{t},\mathcal{P},\mathbf{q}_{t})\\ &-(\mathbf{z}-\mathbf{a}^{*}_{t+1:t+H})\|^{2}\bigg],\end{split}$ (2)
@@
As Figure 1 shows, we apply a hierarchical attention mask to constrain information flow across modules. Action-level states are prevented from attending to trajectory-level or task-level representations, with all tokens causally masked. Specifically, the action state $q_{t}$ cannot attend to the trajectory state $q^{\prime}_{t}$, which enforces a Task → Trajectory → Action hierarchy and avoids information leakage.
III-C Data Collection for Human and Robot
@@
Human Hand Motion Capture System. We capture high-fidelity 3D human hand poses using a calibrated multi-camera system (as shown in Figure 2). Our pipeline uses Google MediaPipe [43] to detect 2D hand keypoints from three views, then fits the parametric MANO [32] 3D hand model to these keypoints by minimizing reprojection error. This yields an accurate 3D hand motion, from which we extract the operational endpoint’s trajectory. The full algorithm is summarized in Algorithm 1. Algorithm details are provided in Appendix A-A.
Algorithm 1 Model-Based 3D Hand Reconstruction
1:
3:
4:Optimized parameters $\boldsymbol{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*}$.
5:
6:procedure ReconstructHandPose($\{\mathbf{k}_{c,j}\},\dots$)
7:  $\boldsymbol{\theta}\leftarrow\mathbf{0},\mathbf{R}\leftarrow\mathbf{I},\mathbf{t}\leftarrow\mathbf{0}$ $\triangleright$ $\triangleright$ Initialize optimizable parameters
8:  $\boldsymbol{\beta}\leftarrow\mathbf{0}$ $\triangleright$ $\triangleright$ Use fixed mean hand shape
9:  $N_{iter}\leftarrow\textsc{IterationScheduler.get\_iterations}$
10:  for $i=1\to N_{iter}$ do
11:   $\{\mathbf{J}_{j}\}\leftarrow\mathcal{M}(\boldsymbol{\beta},\boldsymbol{\theta},\mathbf{R},\mathbf{t})$ $\triangleright$ $\triangleright$ Generate 3D joints via MANO forward pass
13:   Update parameters $\boldsymbol{\theta},\mathbf{R},\mathbf{t}$ via gradient descent on lossloss.
14:  end for
15:  return $\boldsymbol{\theta},\mathbf{R},\mathbf{t}$ as $\boldsymbol{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*}$
16:end procedure
@@
Robot Data Collection. Our platform features a Franka Research 3 arm with a UMI Gripper [10]. An expert teleoperates the robot via a SpaceMouse. We record proprioceptive data and images from a static third-person camera and a wrist-mounted camera. The robot’s end-effector 3D position is directly recorded as its trajectory. Simultaneously, the end-effector’s 3D orientation and the gripper state are also recorded. Combined with the 3D position, these constitute the full action label for policy training. To enhance policy robustness, we randomize the robot’s starting pose for each demonstration, which improves the policy’s ability to recover from errors. For more details, please refer to Appendix A-A.
IV Experiments
Real-World Tasks. As shown in Figure 3, four distinct real-world tasks are designed to evaluate the policy’s capabilities across basic manipulation, language grounding, long-horizon planning, and precise control. All experimental designs in this chapter are centered on two core research objectives: first, to systematically verify the effectiveness of the proposed Traj2Action framework from multiple dimensions including architectural design, component ablation, and cross-task generalization; second, to reveal the key laws of human hand demonstration data in robot policy learning, such as data collection efficiency characteristics, scale effect, and substitution effect for expensive robot teleoperation data. The tasks are detailed as follows:
Task 1: pick up the water bottle (PWB): The robot is tasked with locating a water bottle placed on a tabletop, moving towards it, and grasping it successfully. This task evaluates the model’s fundamental pick-and-place capabilities.
Task 2: pick up the tomato and put it in the yellow/blue tray (PTT): The workspace contains a tomato and two trays, one yellow and one blue. The robot must pick up the tomato and place it into the tray specified by a language command (e.g., “the yellow tray”). This task tests the policy’s ability to task instructions to specific objects and goals.
Task 3: stack the rings on the pillar (SRP): The scene includes a pillar (composed of a yellow column and a blue base), a yellow ring, and a red ring. The robot needs to pick up both rings, one by one, and place them onto the pillar. This task assesses multi-step object manipulation and precision location, which demands that the policy has precise action control.
Task 4: stack the paper cups (SPC): Three paper cups are placed on the table. The robot is required to stack them sequentially to form a single tower. This task evaluates the policy’s ability to handle deformable objects and perform iterative, precise placement.
We categorize these tasks as either Short-Horizon or Long-Horizon. Tasks 1 and 2 are Short-Horizon, evaluating fundamental skills like object grasping and language grounding. In contrast, Tasks 3 and 4 are significantly more demanding Long-Horizon tasks that test the policy’s planning capability. They not only require the aforementioned skills but also critically assess the policy’s ability to perform high-level planning for complex, sequential operations.